# Zhang, Liu, and Zhao

### v4

@@ 9.2 Convergence of ESMM to B-stationarity
The B-stationarity coincides with d-stationarity if $\bar{\Theta}$ is convex. Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (9.2) are B-differentiable. With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (9.2) by $\mathcal{S}^{b}$.
[…]
@@
then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (9.2) is feasible to (9.2) and thus is a B-stationary point of (9.2).
Condition (32) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (9.2) are bounded away from 0. Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (9.2) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (9.2). Nonetheless, it is not easy to verify the technical condition (32) when $G_{\gamma}$ is in a composite structure. The following assumption on $\{\psi_{j}\}_{j\in[J]}$ provides a sufficient condition for (32):
@@
Suppose that Assumption (B7) holds, and that $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set with $0\in\Theta$ and $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Then condition (32) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (9.2) is a B-stationary point of problem (9.2).
Proof 9.7
@@
$\mathcal{A}_{<}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma<0\},\ \mathcal{A}_{>}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma>0\},\text{ and }\mathcal{A}_{=}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma=0\}.$
@@
[…]
@@
Under Assumptions (B1) to (B4) and the same settings of Proposition 9.6, for problem (9.2) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. The following statements hold:
@@
    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (9.2) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (9.2).
@@
For the ESMM algorithm with exact penalization, the limit point of generated sequence is a composite strong d-stationarity to (9.2) with probability 1 according to Theorem 9.8, and thus is a global minimizer of (9.2) when $\varepsilon$ is sufficiently large and $\eta$ is close to 0, by the equivalence between composite strong d-stationarity and global optimality in Proposition 8.2 with appropriate assumptions when needed. Furthermore, such a global minimizer of the penalized problem (9.2) should be also a global minimizer of (9.2) since it is feasible according to Proposition 9.6. Under such settings, the asymptotic consistency of the PADR framework for the constrained case in Theorem 9.3 turns into asymptotic optimality.
@@ 9.3 Experiments on constrained newsvendor problem
In this subsection, we compare PADR with prescriptive methods and the penalized RKHS-DR method from Bertsimas and Koduri (2021, Section 6.2).
@@
Since DR-based methods may yield infeasible decisions, in Figure 1u we report the test cost of the order decisions, projected onto the convex feasible region when necessary, along with the feasible frequency of the original decisions provided by PADR and RKHS-DR.
@@
(s) Basic results
(t) Sparse demand model
(u) Dense demand model
@@
With the basic demand model, PADR converges to SIMOPT and attains 100% feasibility when $n\geq 200$. In comparison, RKHS-DR and prescriptive methods display notable gaps to SIMOPT, with the feasible frequency of RKHS-DR being only around 80%. Figures 1t and 1u show the numerical results with varying feature dimensions. Similar to the unconstrained cases, PADR outperforms the other methods in the sparse case when $p<20$ and in the dense case when $p<50$. Notably, with penalization, PADR achieve 100% feasibility when $p<20$. Although PADR can not guarantee 100% feasibility of the decision compared to prescriptive methods, it still provides satisfactory performance via projection.
@@
We next consider a capacity constraint $C(z_{1})+C(z_{2})\leq C_{0}$ with the concave capacity cost function (18) for the two products. We set $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and $C_{0}=50$ with the same probability for the constraint being active as the experiment in Section 9.3.1. Since the projection of infeasible decisions to the nonconvex feasible region may be far away to the optimal solutions, we only report the average test cost of feasible decisions in Figure 1x, with a careful hyperparameter selection to ensure high feasible frequencies.
We use the standard MM algorithm to solve the penalized ERM problem of RKHS-DR, referred to here as RKHS-DR(MM). For prescriptive methods and SIMOPT, we solve the two-dimensional nonconvex-constrained problems using grid search. The sof method in Kallus and Mao (2023)is not suitable for this nonconvex-constrained problem, and it may require the study of the second-order perturbation analysis for stochastic optimization, which is beyond the scope of this paper.
For the problem with the basic demand model, both PADR and RKHS-DR attain 100% feasibility on the test set. PADR converges to SIMOPT with significantly fewer training samples than other benchmarks. Figures 1w and 1x show the results across varying feature dimensions. PADR outperforms the other methods in both sparse and dense cases when $p\leq 30$, achieving 90 % $\sim$ 100% feasibility in the sparse case when $p\leq 20$ and in the dense case when $p\leq 30$. Under the dense demand model, RKHS-DR performs well in low-dimensional cases, but its performance deteriorates as $p$ increases. Prescriptive methods—kNN (in the dense case), cart, and rf—maintain a stable yet noticeable gap to SIMOPT, which is possibly due to their limitations to approximate conditional distributions with the sample set that is not sufficiently large. The relatively small gap of PADR to SIMOPT shows PADR’s capability to learn the underlying decision rule structure in problems with nonconvex constraints while maintaining a high probability of feasibility when $p\leq 30$.
@@
(v) Basic results
(w) Sparse demand model
(x) Dense demand model

### v5

@@ 9.2 Convergence of ESMM to B-stationarity
The constrained ERM problem (33)can be reformulated as follows with $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $\Xi$:
@@
[…]
Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ are Bouligand-differentiable (B-differentiable), i.e., they are locally Lipschitz continuous and directionally differentiable on $\mathbb{R}^{d}$, and that $\Theta$ is a compact convex set. Since $G_{\gamma}$ is nonconvex, we are interested in computing Bouligand-stationary (B-stationary) points of (36), the defining condition of which is an extension of d-stationarity to nonconvex-constrained optimization problems. With $\mathcal{I}>0$, we construct the following penalized problem
$\underset{\displaystyle\begin{subarray}{c}\theta\in\Theta\end{subarray}}{\mathrm{min}}\quad V(\theta;\lambda)\coloneqq F(\theta)+\lambda G_{\gamma}(\theta).\hfil\hfil\hfil\hfil$ (37)
According to Cui and Pang (2021a, Proposition 9.2.2), problem (33) with some regularity conditions has the exact penalty property such that with a finite penalty parameter $\lambda$, any d-stationary point of (37) is a B-stationary point of (36). Since $\max\{\cdot,0\}$ is nondecreasing and convex, it is clear that the upper surrogate functions of $G_{\gamma}(\cdot)$ can be constructed by plugging surrogates of $\{\psi_{j}(f(\cdot,\xi))\}_{j\in[J]}$ into $\max\{\cdot,0\}$, maintaining the corresponding surrogation properties. Therefore, with the exact penalty property and Assumptions (B1)-(B6) imposed in the context of problem (37), the ESMM algorithm is ready to solve the penalized problem (37) with convergence to B-stationary points of (36). According to (Cui and Pang 2021b), for a constrained problem with a B-differentiable objective, a B-stationary point is defined as follows.
@@
The B-stationarity coincides with d-stationarity if $\bar{\Theta}$ is convex. Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (36) are B-differentiable. With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (36) by $\mathcal{S}^{b}$.
With Assumptions (B1)-(B6) imposed in the context of the penalized problem (37), we can apply the ESMM algorithm to solve (37) and obtain the same convergence results in terms of d-stationarity as in Theorems 3.7-8.12. To further obtain the convergence to B-stationarity, we intend to connect d-stationarity of (37) with B-stationarity of (36) by applying the following exact penalty result from Cui and Pang (2021b). Note that by the finiteness of $\Xi$ and compactness of $\Theta$, for any $\xi\in\Xi$, $f(\cdot,\xi)$ is Lipschitz continuous with modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$ and uniformly bounded with $\left\lVert f(\cdot,\xi)\right\rVert_{\infty}\leq M_{f}\coloneqq L_{f}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, and $\varphi(\cdot,\xi)$ is Lipschitz continuous on $M_{f},M_{f}]^{d}$ with uniform modulus denoted by $L^{\prime}_{\varphi}$.
Lemma 9.7 (Cui and Pang (2021b), Proposition 9.2.2)
@@
then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (37) is feasible to (36) and thus is a B-stationary point of (36).
Condition (38) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (36) are bounded away from 0. Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (37) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (37). Nonetheless, it is not easy to verify the technical condition (38) when $G_{\gamma}$ is in a composite structure. The following assumption on $\{\psi_{j}\}_{j\in[J]}$ provides a sufficient condition for (38):
@@
Suppose that Assumption (B7) holds, and that $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set with $0\in\Theta$ and $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Then condition (38) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (37) is a B-stationary point of problem (36).
Proof 9.9
@@
[…]
@@
$G_{\gamma}(\cdot,\xi;j)^{\prime}(\bar{\theta};\bar{v})=\max\big\{\psi_{j}^{\prime}(f(\bar{\theta},\xi);f(\cdot,\xi)^{\prime}(\bar{\theta};\bar{v})),\,0\big\}=\max\left\{\psi_{j}^{\prime}\left(\bar{z};-\frac{\bar{z}}{\left\lVert\bar{\theta}\right\rVert}\right),0\right\}=0.$
@@
Under Assumptions (B1) to (B4) and the same settings of Proposition 9.8, for problem (37) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. The following statements hold:
@@
    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (37) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (36).
@@
For the ESMM algorithm with exact penalization, the limit point of generated sequence is a composite strong d-stationarity to (37) with probability 1 according to Theorem 9.10, and thus is a global minimizer of (37) when $\varepsilon$ is sufficiently large and $\eta$ is close to 0, by the equivalence between composite strong d-stationarity and global optimality in Proposition 8.2 with appropriate assumptions when needed. Furthermore, such a global minimizer of the penalized problem (37) should be also a global minimizer of (36) since it is feasible according to Proposition 9.8. Under such settings, the asymptotic consistency of the PADR framework for the constrained case in Theorem 9.5 turns into asymptotic optimality.
@@ 9.3 Experiments on constrained newsvendor problem
In this subsection, we implement the prescriptive methods, which ensure feasibility by solving the individual constrained problem for each sample. Additionally, we implement NN-DR equipped with $\ell_{1}$-penalization and RKHS-DR methods following the penalized formulation proposed by Bertsimas and Koduri (2021, Section 6.2).
@@
Since DR-based methods (i.e., PADR, NN-DR, and RKHS-DR) may yield infeasible decisions, in Figure 1aa we report the test cost of the order decisions, projected onto the convex feasible region when necessary, along with the feasible frequency of their original decisions.
@@
(y) Basic results
(z) Sparse demand model
($a$) Dense demand model
@@
Figure 1y shows that with the basic demand model, PADR approaches SIMOPT and achieves 100% feasibility when $n\geq 200$. In contrast, other benchmarks exhibit noticeable performance gaps relative to SIMOPT. Particularly, RKHS-DR achieves roughly 80% feasibility and NN-DR exhibits unstable feasibility performance. Figures 1z and 1aa present the results under varying feature dimensions. Consistent with the unconstrained setting, PADR outperforms other methods in the sparse model when $p<20$ and in the dense model when $p<50$. Notably, by the utilization of $\ell_{1}$ penalization, PADR achieves 100% feasibility when $p<20$. Moreover, in the dense demand model, PADR exhibits the clear edge over prescriptive methods by achieving better empirical performance on test costs, while it fails to be feasible at 10% - 20% probability.
@@
We next consider a capacity constraint $C(z_{1})+C(z_{2})\leq C_{0}$ with the concave capacity cost function (18) for the two products. We set $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and $C_{0}=50$ with the same probability for the constraint being active as the experiment in Section 9.3.1. Since the projection of infeasible decisions to the nonconvex feasible region may be far away to the optimal solutions, we only report the average test cost of feasible decisions in Figure 1ad, with a careful hyperparameter selection to ensure high feasible frequencies.
We use the standard MM algorithm to solve the penalized ERM problem of RKHS-DR, referred to here as RKHS-DR(MM). For prescriptive methods and SIMOPT, we solve the two-dimensional nonconvex-constrained problems using grid search. The SOF method in Kallus and Mao (2023)is not suitable for this nonconvex-constrained problem, and it may require the study of the second-order perturbation analysis for stochastic optimization, which is beyond the scope of this paper. For the other prescriptive methods, we solve two-dimensional nonconvex constrained problems by grid search.
For the problem with the basic demand model, both PADR and RKHS-DR attain 100% feasibility on the test set. PADR converges to SIMOPT with significantly fewer training samples than other benchmarks. Figures 1ac and 1ad show the results across varying feature dimensions. PADR outperforms the other methods in both sparse and dense cases when $p\leq 30$, achieving 90 % $\sim$ 100% feasibility in the sparse case when $p\leq 20$ and in the dense case when $p\leq 30$. Under the dense demand model, RKHS-DR performs well in low-dimensional cases, but its performance deteriorates as $p$ increases. Prescriptive methods—kNN (in the dense case), CART, and RF—maintain a stable yet noticeable gap to SIMOPT, which is possibly due to their limitations to approximate conditional distributions with the sample set that is not sufficiently large. The relatively small gap of PADR to SIMOPT shows PADR’s capability to learn the underlying decision rule structure in problems with nonconvex constraints while maintaining a high probability of feasibility when $p\leq 30$. NN-DR obtains relatively lower test costs than PADR and RKHS-DR; however, its feasibility rate is substantially lower than those of the other two methods. This suggests that the decision rule yielded by NN-DR may face difficulties in effectively addressing nonconvex-constrained optimization problems.
@@
(ab) Basic results
(ac) Sparse demand model
(ad) Dense demand model
@@
Thanks to the Associate Editor who suggested a hybrid idea during the review process to address the feasibility issue of DR-based methods, we implement a hybrid strategy by combining the decision rule with an auxiliary method that ensures feasibility (such as RF). Specifically, for the Newsvendor problem with a nonconvex constraint, we combine PADR and NN-DR with the prescriptive method RF, referred to as PADR(hybrid) and NN-DR(hybrid). Each method applies the corresponding decision rule first; if the output is infeasible, the RF solution is used as a fallback. We compare the performance of these hybrid methods with pure RF and CART methods, as shown in Figure 1af. Both PADR(hybrid) and NN-DR(hybrid) consistently outperform RF, with the exception when $p=50$. Under the sparse demand model, NN-DR(hybrid) outperforms PADR(hybrid) and shows greater improvements over RF. Under the dense demand model, the performance of both hybrid methods is close to SIMOPT, with PADR(hybrid) slightly better. Overall, these preliminary results demonstrate the promise of hybrid approaches in improving both feasibility and performance. This is especially beneficial for contextual SP problems with complex, nonconvex constraints, and we believe this avenue deserves further investigation.
@@
(ae) Sparse demand model
(af) Dense demand model
Newsvendor problem with a nonconvex constraint: hybrid approaches
10 Examples Satisfying or Violating Assumption (A4)
In this section, we present several examples in which the SP problems either satisfy or violate Assumption (A4). We conclude with a brief discussion on a potential analytical framework to establish sufficient conditions for the Lipschitz continuity of the optimal decision rule.
10.1 Problems satisfying Assumption (A4)
A canonical class of problems ensuring Assumption (A4) is quantile regression, such as the Newsvendor problem, which in fact served as the motivation for the Lipschitz continuity assumption in our paper. Additionally, as extensively discussed in Shapiro et al. (2021a, Chapter 6), risk-averse optimization also yields quantile-based solutions. For example, Average Value-at-Risk (AV@R) attains its minimum at the Value-at-Risk (V@R), which by definition is a quantile of the underlying distribution (see Section 6.2.4). Section 6.4.2 in the above reference further shows that in the risk-averse Newsvendor problem, where the objective is a composition of a risk measure and the inventory cost, if the risk measure is real-valued and coherent (e.g., AV@R or the mean-upper-semideviation of order $p$ in Section 6.3.2), the optimal order quantity remains a quantile of a distribution that depends on the risk measure and the inventory parameters. In these situations, determining $z^{*}$ reduces to estimating (or “regressing”) the conditional quantile function based on $X$. Thus, the Lipschitz continuity of $z^{*}$ hinges on that of the conditional quantile function. Under the additive model $Y(x)=\bar{Y}(x)+\epsilon$, where $\epsilon$ is an independent and zero-mean random variable, and $X$ is compact, the continuity of the mean function $\bar{Y}(x)$ implies the desired Lipschitz continuity.
Another class of problems satisfying Assumption (A4) is a two-stage stochastic optimization problem:
$\min_{z\in\mathcal{Z}}\quad g(z)+\mathbb{E}\left[Q(Az,Y)\mid X=x\right],$
where
$Q(u,Y)=\min\{q^{\top}v:Wv=Y-u,v\geq 0\}.$
Here, $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is convex, $\mathcal{Z}$ is a closed convex set, and $\mathbb{P}_{Y|X=x}$ is a Borel probability measure with finite first moment. Under assumptions of complete recourse and dual feasibility for the second-stage problem, Römisch and Schultz (1993, Theorem 2.2) show that if the function $\mathbb{E}\left[Q(\cdot,Y)\mid X=x\right]$ is strongly convex for all $x\in\mathcal{X}$, then the optimal solution set is a singleton, and there exist constants $L>0$ and $\delta>0$ such that
$\left\lVert Af^{*}(x_{1})-Af^{*}(x_{2})\right\rVert\leq L\mathcal{W}_{1}(\mathbb{P}_{Y\mid x_{1}},\mathbb{P}_{Y\mid x_{2}})^{1/2},$
whenever $\mathcal{W}_{1}(\mathbb{P}_{Y\mid x_{1}},\mathbb{P}_{Y\mid x_{2}})\leq\delta$, where $\mathcal{W}_{1}$ denotes the $L_{1}$-Wasserstein distance. Consequently, if $A$ is invertible, $\mathcal{X}$ is compact, and $Y(x)=\bar{Y}(x)+\epsilon$ with continuity of $\bar{Y}(\cdot)$, the Lipschitz continuity of the optimal policy follows. A subsequent work by Römisch and Schultz (1996) further investigates the Lipschitz stability of the optimal solution sets with respect to a pseudo-metric over the space of probability measures, which may facilitate the establishment of Lipschitz continuous solution mappings in problems with non-unique solutions.
10.2 Problems not satisfying Assumption (A4)
In this part, we present examples where Assumption (A4) does not hold. Consider the additive random variable model $Y(x)=\bar{Y}(x)+\epsilon$ with a discontinuous mean function $\bar{Y}(x)$ on $\mathcal{X}$. Then the resulting optimal decision rule will be discontinuous. A notable instance arises in regression problems with the mean squared error loss as the cost function, where the optimal decision rule coincides with $\bar{Y}(x)$ under normally distributed noise $\epsilon$. If $\bar{Y}(x)$ is discontinuous or discrete, e.g., due to constraints that depend on $x$, the corresponding optimal decision rule will also be discontinuous.
For many problems, the set of optimal decisions is inherently disconnected across scenarios. One case is when the feasible region $\mathcal{Z}$ itself is disconnected, so that the optimal solutions conditional on different features may lie in different connected components. Another case is when the problem structure yields discrete optimal solutions, such as when $\mathcal{Z}$ is a polytope and the optimal decision always lies at an extreme point. In these situations, small perturbations of the input can cause abrupt jumps in the optimal decision rule, thereby not satisfying Assumption (A4). Addressing such cases calls for the development of Heaviside function classes (Fang et al. 2025) including decision trees to capture the discontinuous components of the decision rule. Advancing such methods constitutes a promising and important direction for future research.
To conclude this section, we briefly outline a potential direction for analyzing sufficient conditions for the Lipschitz continuity of the optimal decision rule. A principled approach is to first characterize conditions on the function $\Phi(z,x)\coloneqq\mathbb{E}_{Y}[\varphi(z,Y)\mid X=x]$ (defined over $\mathcal{Z}\times\mathcal{X}$) that ensure Lipschitz continuity of the solution mapping $z^{*}(x)\in\arg\min_{z\in\mathcal{Z}}\Phi(z,x)$, using implicit function theorems for variational inequalities. In particular, with the function $\Phi(z,x)$ being continuously differentiable in $z$ and the generalized equation $0\in\nabla_{z}\Phi(z,x)+\mathcal{N}_{\mathcal{Z}}(z)$ being sufficient for optimality, Dontchev and Rockafellar (2009) provide conditions on the mapping $x\mapsto\nabla_{z}\Phi(\bar{z},x)$ under which the mapping $z^{*}(x)$ is locally Lipschitz continuous (e.g., Dontchev and Rockafellar (2009, Theorem 2B.5)). It thus reduces to identify structural properties of $\varphi(z,Y)$ and conditional distributional properties of $Y$ such that those regularity conditions hold, such as growth properties of $\Phi$. While we recognize that verifying these conditions is nontrivial in general, this framework offers a promising direction for theoretically analyzing the regularity of optimal policies in contextual SP, which deserves further exploration.
11 Surrogation Construction for Problems in Section 4
In this section we provide the specific surrogation construction of $F(\cdot,\xi^{s})$ with $\xi^{s}\in\Xi$ for problems in Section 4, including the Newsvendor problem, the Newsvendor problem with additional concave cost, and the product placement problem.
For the Newsvendor problem, the cost function $\varphi(z,\xi^{s})=c_{b}\max\{y^{s}-z,0\}+c_{h}\max\{z-y^{s},0\}$ is the sum of two monotonic convex functions. By utilizing surrogation construction for Case 2, the surrogate function of $F(\theta,\xi^{s})=\varphi(f(\theta,\xi^{s}),\xi^{s})$ at $\theta^{\nu}$ is constructed as follows,
$\widehat{F}(\theta,\xi^{s};\theta^{\nu},I)=c_{b}\max\left\{\widehat{f}(\theta,\xi^{s};I_{2})-y^{s},0\right\}+c_{h}\max\left\{y^{s}-\widecheck{f}(\theta,\xi^{s};I_{1}),0\right\},$
where $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\nu},\xi^{s})$ and $\widehat{f}(\theta,\xi^{s};I_{2})$ and $\widecheck{f}(\theta,\xi^{s};I_{1})$ are the upper and lower surrogate functions of $f(\theta,\xi^{s})$ respectively.
For the Newsvendor problem with the additional concave piecewise-affine cost function, we denote the concave cost function by the general form $C(z)\coloneqq\min_{i\in[K_{C}]}\{a_{i}z+b_{i}\}$. In this case, the cost function is
$\varphi(z,\xi^{s})$ $=c_{b}\max\{y^{s}-z,0\}+c_{h}\max\{z-y^{s},0\}+\min_{i\in[K_{C}]}\{a_{i}z+b_{i}\}$
$=c_{b}\max\{y^{s}-z,0\}+c_{h}\max\{z-y^{s},0\}-\max_{i\in[K_{C}]}\{-a_{i}z-b_{i}\},$
which satisfies Case 1. Following the construction for Case 1, the surrogate function of $F(\theta,\xi^{s})=\varphi(f(\theta,\xi^{s}),\xi^{s})$ at $\theta^{\nu}$ is constructed as below,
$\widehat{F}(\theta,\xi^{s};\theta^{\nu},I)=$ $c_{b}\max\left\{\widehat{f}(\theta,\xi^{s};I_{2})-y^{s},0\right\}+c_{h}\max\left\{y^{s}-\widecheck{f}(\theta,\xi^{s};I_{1}),0\right\}$
$-\left(\max\{(-a_{I_{3}}),0\}^{\top}\widecheck{f}(\theta,\xi^{s};I_{1})+\min\{(-a_{I_{3}}),0\}^{\top}\widehat{f}(\theta,\xi^{s};I_{2})-b_{I_{3}}\right),$
where $I_{1}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\nu},\xi^{s})$, $I_{2}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\nu},\xi^{s})$, and $I_{3}\in\mathcal{I}^{\varepsilon}_{C}(f(\theta^{\nu},\xi^{s}),\xi^{s})$. Consistent with the surrogation for Case 1, the surrogation index $I=(I_{1},I_{2},I_{3})$ in this case includes the surrogation index for the outer function $C(\cdot)$ at the point $f(\theta^{\nu},\xi^{s})$.
Finally, for the product placement problem, we first convert the two-stage problem into a single-stage convex problem by extending the auxiliary variable $f$:
$\min_{\begin{subarray}{c}z\geq 0\\ v^{s}\geq 0,\ s\in[n]\end{subarray}}\quad\frac{1}{n}\sum_{s=1}^{n}\left\{c^{\top}z+g^{\top}v^{s}+b^{\top}\max\{y^{s}+Wv^{s}-z,0\}\right\}.$
Since $c\geq 0$ and $b\geq 0$, for the component $\varphi(z,v,\xi^{s})=c^{\top}z+g^{\top}v+b^{\top}\max\{y^{s}+Wv-z,0\}$, we construct the surrogate function of $F(\theta,v,\xi^{s})\coloneqq\varphi(f(\theta,\xi^{s}),v,\xi^{s})$ at $(\theta^{\nu})$ as
$\widehat{F}(\theta,v,\xi^{s};\theta^{\nu},I)=c^{\top}\widehat{f}(\theta,\xi^{s},I_{2})+g^{\top}v+b^{\top}\max\left\{y^{s}+Wv-\widecheck{f}(\theta,\xi^{s};I_{1}),0\right\},$
where $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\nu},\xi^{s})$.
12 Additional Experimental Results
In this section, we further conduct experiments on the Newsvendor problem using alternative demand models that share a fixed linear component but with different nonlinear structures, including quadratic, cubic, piecewise-affine and sine functions. These experiments are designed to assess the robustness of PADR and benchmark methods across varying nonlinear models. By incorporating an explicit linear term in all models, we aim to examine whether the nonlinear effects still manifest when combined with a linear component.
Specifically, we consider the following demand models:
$\bar{Y}_{\text{ec1}}(X)=3(X_{1}+X_{2})+\omega\cdot\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+20,$ (39)
$\bar{Y}_{\text{ec2}}(X)=3(X_{1}+X_{2})+\omega\cdot 5(X_{1}-X_{2})^{2}+10,$ (40)
$\bar{Y}_{\text{ec3}}(X)=3(X_{1}+X_{2})+\omega\cdot[5(X_{1}-X_{2})^{3}-15(X_{1}-X_{2})]+50,$ (41)
$\bar{Y}_{\text{ec4}}(X)=3(X_{1}+X_{2})+\omega\cdot[4\sin(\pi X_{1})+\max\{16X_{2},-20X_{2}\}]+30.$ (42)
We report the results under the above models in Figures 1ah-1an.